# Effective Heterogeneous Federated Learning via Efficient Hypernetwork-based Weight Generation

### v1

@@ 1. Introduction
Previous solutions targeting device heterogeneity in federated learning have leveraged techniques such as knowledge distillation (Li and Wang, 2019; He et al., 2020), local model pruning (Li et al., 2021a, 2020b), and adaptive model scaling (Ilhan et al., 2023; Diao et al., 2020). In knowledge distillation, the server is required to maintain a well-balanced public dataset, which is not always easy to secure. Local model pruning relaxes such constraints but requires full model training at clients for model compression. Finally, while adaptive model splitting can be a viable approach, we later show that the heterogeneity of clients leads to information imbalance when model parameters are gathered at the server (Shen et al., 2024; Yao et al., 2021).
@@
With heterogeneous platforms utilizing multi-exit architectures, all clients execute the model up to the first exit, while only a subset of powerful clients proceed to the final. This results in varying sample availability (i.e., information disparity) across different exits at the server, particularly affecting deeper layers with fewer samples for aggregation. Our approach mitigates this challenge using hypernetworks to infer and generate weights for layers lacking sufficient samples, enhancing model aggregation effectiveness. However, despite hypernetworks operating on the server, their resource requirements are extremely high; resulting in large memory usage and long computation times. To address this, we propose a Low-Rank Factorization (LRF)-based compression scheme for neural network parameters, aimed at optimizing hypernetwork operations for federated learning (Sec. 4.3).
We integrate these techniques as a heterogeneous federated learning framework, HypeMeFed, and present extensive evaluations via GPU-based emulations and a real-world testbed of heterogeneous embedded platforms. HypeMeFed outperforms all state-of-the-art baselines and our testbed results show that it improves the model accuracy by 5.13% compared to FedAvg. Furthermore, with its LRF-based optimization, HypeMeFed reduces the hypernetwork’s memory requirement by 98.22% and achieves 1.86 faster hypernetwork operations compared to a naive approach. Overall, our results suggest that HypeMeFed is an effective solution to enable federated learning for heterogeneous clients.
@@
We identify practical challenges in federated learning for heterogeneous clients and propose HypeMeFed. HypeMeFed integrates a multi-exit neural network architecture with hypernetwork-based model weight generation to offer a practical solution for supporting federated learning on systems consisting of heterogeneous clients.
While hypernetworks offer an environment to generate model weights for portions of the network that lack client-sourced parameters, their naive use results in significant computational and memory overhead. We propose an LRF-based approach to compress and optimize its operations to a practically applicable scale, making hypernetworks suitable for federated learning scenarios.
@@ 2. Related work and Background
We begin by positioning our work within the current literature and discuss the foundational concepts and methodologies relevant to our study.
Device Heterogeneity in Federated Learning. While federated learning research has mostly focused on homogeneous models with little consideration of client-side computing capabilities (Diao et al., 2020), supporting federated learning for heterogeneous clients is crucial. It allows for tailored models that adapt to diverse computational capabilities, data distributions, and privacy requirements of various devices (Zhang et al., 2022; Bonawitz et al., 2019; Li et al., 2020a). Unlike homogeneous approaches, heterogeneous federated learning leverages each client’s unique capabilities, ensuring that even resource-constrained devices contribute effectively to enhance model robustness, maximize the potential of collective data, and lead to better generalization across tasks and environments (Lin et al., 2020).
One early approach to achieve device heterogeneity in federated learning is federated distillation (FD) (Li and Wang, 2019; Seo et al., 2022), which leverages knowledge distillation (Hinton et al., 2015). Instead of sharing model weights, FD aggregates class scores from a large public dataset computed locally at each client. FD accommodates diverse client models but relies heavily on the quality and availability of public datasets and requires this data for all clients, posing challenges for resource-limited platforms.
An alternative approach involves pruning local models based on the Lottery Ticket hypothesis (Frankle and Carbin, 2018). In LotteryFL (Li et al., 2020b) and Hermes (Li et al., 2021a), each client learns a subnetwork of the global model (i.e., lottery ticket network) through iterative pruning, rewinding, and training, exchanging only sparse pruned parameters. However, this iterative process is practically challenging for resource-limited clients, as they must initially train the heavy full model to determine which weights to prune. This requires a large number of rounds for model convergence and can cause training delays.
@@
Multi-Exit Network Architecture. As the name implies, multi-exit architectures, inject multiple (early) exit points within a neural network to enhance inference latency by enabling adaptive inference and eliminating unnecessary operations (Laskaridis et al., 2021; Lee et al., 2024a). Teerapittayanon et al. introduced BranchNet, a pioneering multi-exit deep neural network featuring multiple classifiers integrated within the baseline model to improve computational latency (Teerapittayanon et al., 2016). Given that mobile and embedded devices often experience lagging inference due to resource limitations, multi-exit architectures have been often applied in various real-world sensing applications within mobile and embedded computing environments (Kouris et al., 2022; Leontiadis et al., 2021; Kim et al., 2020).
Hypernetworks. Hypernetworks are neural networks designed to generate weights for another neural network (Ha et al., 2016), offering flexibility in applications like computer vision (Jia et al., 2016; Klocek et al., 2019) and 3D scene representation (Littwin and Wolf, 2019; Sitzmann et al., 2020). In federated learning, hypernetworks are used to enhance model initialization and personalization (Shamsian et al., 2021) and to determine aggregation ratios for model personalization (Ma et al., 2022).
@@ 3.2. Preliminary Study
In Figure 3 (a) we also plot four additional cases where our proposed solutions in Section 3.1 were applied in different combinations. As we allocate clients with their largest-adoptable model, we text cases where neither the multi-exit approach nor the hypernetwork is used (“ME X/HN X”), cases where one of the two is used (“ME O/HN X” and “ME X/HN O”), and a final case there both approaches are integrated into the federated learning process (“ME O/HN O”). We use these plots to examine the feasibility of applying our solutions for federated learning scenarios. The results in Figure 3 (a) suggest that in the absence of the multi-exit architecture, convergence issues arise due to misaligned feature spaces, while the system faces issues due to per-layer information disparity when without the hypernetwork-based weight generations. As a result, the case with both multi-exit and hypernetwork (“ME O/HN O”) shows the highest performance, close to that of an ideal case where all clients exploit the full model (“Large (All)”). These results emphasize the need for both the multi-exit and hypernetwork approaches.
@@
Without multi-exit layers, the one-layer (S) and two-layer (M) models differ significantly from the full model (L), with CKA values of 0.6761 and 0.6764, respectively. However, incorporating multi-exit architecture improves feature alignment noticeably, yielding CKA values of 0.7812 and 0.9407 for these configurations. Comparisons like “M-L (X-O)” and “M-L (O-X)” in Figure 3 (b) show considerable differences in feature alignment depending on the presence of multi-exit layers, with CKA values of 0.7165 and 0.8343, respectively. These findings highlight the effectiveness of multi-exit architectures in depth-wise split networks for federated learning, enhancing feature space alignment and potentially boosting overall model performance.
@@
Overall, our preliminary results suggest the potential effectiveness of leveraging the multi-exit architecture and hypernetworks for heterogeneous federated learning. Using these results as empirical evidence, we discuss the details of our proposed HypeMeFed design in the following section.
@@ 4.1. Overview of HypeMeFed
The server distributes global model parameters to clients based on their computing resources. The full global model is configured with multiple exit layers organized in a depth-wise manner ( ). Clients with higher computational resources receive the full model, while the ones that are limited receive a smaller subset of the model split at earlier exit layers ( ). (c.f., Sec. 4.2).
@@ 4.2. Multi-Exit Neural Network Architecture
The placement and number of exit layers in HypeMeFed are key design choices. In this work, we use three exit layers positioned at one-third, two-thirds, and the end of the federated learning model to exploit a range of features that the model embeds (Park et al., 2023). This setup allows the categorization of clients into three groups based on their computational capabilities: the first group, with limited resources, uses a subnetwork up to the first exit layer; the second group, with moderate resources, uses up to the second exit; and the third group, with ample resources, uses the full network with all three exits. During local training, clients compute outputs for all exit layers included in their allocated model. For instance, a client with sufficient resources generates predictions from all three exit layers and a resource limited client will compute only up to the first. Each layer’s output contributes to the overall loss, summing the losses from all exits—a standard practice in multi-exit neural networks (Laskaridis et al., 2021; Teerapittayanon et al., 2016; Lee et al., 2024a).
@@
HypeMeFed facilitates seamless aggregation of trained model parameters despite heterogeneous network architectures by training and sharing all exit layer parameters. Specifically, HypeMeFed can easily be integrated with various federated learning schemes like FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020a), and SCAFFOLD (Karimireddy et al., 2020) without modifying the local training and aggregation process. In this work, HypeMeFed adopts the FedAvg approach, a straightforward method for averaging model parameters during aggregation.
@@ 4.3. Efficient Hypernetwork Design for Multi-exit Architectures
To integrate hypernetworks in federated learning, a straightforward approach involves vectorizing parameters from preceding layers and feeding them into a multi-layer perceptron (MLP) model. The central server, which aggregates parameters from clients with different computational capabilities, possesses information on both preceding and subsequent layer parameters. This allows the server to utilize parameter information from clients computing up to the second and third exits to train the hypernetwork and understand relationships between adjacent exit parameters. However, neural networks typically entail numerous parameters, demanding substantial memory and time for hypernetwork training. Despite HypeMeFed’s hypernetworks operating on a resource-rich server, a naïve hypernetwork design can be computationally intensive and memory-consuming, posing practical challenges in real-world deployments. To tackle this, we propose a more efficient hypernetwork architecture that balances effective parameter generation with the computational constraints of federated learning environments.
HypeMeFed introduces an efficient hypernetwork architecture using a low-rank factorization (LRF) strategy. Instead of vectorizing and reconstructing all model parameters, which is computationally intensive, LRF compresses essential neural network parameters, reducing the computational load and memory requirements. By leveraging the inherent sparsity in neural networks, similar to pruning techniques, this approach is both feasible and efficient (He et al., 2017; Frankle and Carbin, 2018; Evci et al., 2022).
@@ 4.4. Hypernetwork Implementations
Given the early-exit architecture of HypeMeFed, which splits the model into three subnetworks, we employ two hypernetworks: one to predict the second exit parameters from the first exit parameters () and another to predict the third exit parameters from the second (). These hypernetworks are trained at the server using parameter samples collected from clients each round. Clients executing the model up to the second exit provide samples for , while those executing the full model provide samples for .
@@
HypeMeFed selectively generates only the core parameters of the feature extractor, specifically the convolutional layers, while excluding the batch normalization (BN) and classification layers. This approach is based on two key empirical observations: (i) BN and classification layers do not significantly enhance global model performance, and (ii) since these layers are closely related to local data, globally generating them can dilute the unique knowledge of each client, compromising model personalization. We demonstrate this in Section 5.3, where we evaluate personalization accuracy after a few epochs of local fine-tuning of the global model with client data.
Once client model parameter data is collected at each federated learning round and the hypernetworks are trained using the operations described above, HypeMeFed uses the fine-tuned hypernetworks to predict subsequent layer parameters based on the compressed representations. This approach ensures that clients with varying computational capacities can fairly participate in the federated learning process by generating the necessary weight information for effective model convergence and improved performance.

### v2

@@ 1. Introduction
Existing solutions for device heterogeneity in federated learning include techniques like knowledge distillation (Li and Wang, 2019; He et al., 2020), local model pruning (Li et al., 2021a, 2020b), and adaptive model scaling (Ilhan et al., 2023; Diao et al., 2020). Knowledge distillation requires a well-balanced public dataset for transferring knowledge from larger models, which is often impractical in privacy-sensitive settings. Local model pruning reduces resource demands, but still requires full-model training in the initial phase, which is infeasible for many resource-limited devices. Adaptive model scaling, either width-wise or depth-wise, adjusts model sizes based on client resource availablity, but faces the issue of information imbalance during server aggregation, where smaller models provide less informative updates (Shen et al., 2024; Yao et al., 2021). These limitations highlight the need for approaches that balance resource optimization with effective neural network parameter aggregation in heterogeneous environments.
@@
With heterogeneous platforms utilizing multi-exit architectures, all clients execute the model up to the first exit, while only a subset of powerful clients proceed to the final. This results in varying sample availability (i.e., information disparity) across different exits at the server, particularly affecting deeper layers with fewer samples for aggregation. Our approach mitigates this challenge using hypernetworks to infer and generate weights for layers lacking sufficient samples, enhancing model aggregation effectiveness. However, despite hypernetworks operating on the server, their resource requirements are extremely high; resulting in large memory usage and long computation times. We emphasize that hypernetworks can struggle to run effectively even on powerful server platforms due to their memory requirements (Ivison et al., 2022; Chauhan et al., 2024). To address this, we propose a Low-Rank Factorization (LRF)-based compression scheme for neural network parameters, aimed at optimizing hypernetwork operations for federated learning (Sec. 4.3). This LRF-based approach ensures that hypernetworks can generate parameters efficiently at the server, making HypeMeFed scalable and deployable in real-world heterogeneous settings.
We integrate these techniques as a heterogeneous federated learning framework, HypeMeFed, and present extensive evaluations via GPU-based emulations and a real-world testbed of heterogeneous embedded platforms. HypeMeFed outperforms all state-of-the-art baselines, and our testbed results show that it improves model accuracy by 5.13% compared to FedAvg. Furthermore, with its LRF-based optimization, HypeMeFed reduces the hypernetwork’s memory requirement by 98.22% and achieves 1.86 faster hypernetwork operations compared to a naive approach. Overall, our results suggest that HypeMeFed is an effective solution to enable federated learning for heterogeneous clients. We provide an open-sourced implementation of HypeMeFed at https://github.com/eis-lab/HypeMeFed.
@@
We identify practical challenges in federated learning for heterogeneous clients and propose HypeMeFed, a novel framework that combines a multi-exit neural network architecture with hypernetwork-based model weight generation. This approach supports effective federated learning across clients with diverse computational capabilities by addressing the issue of layer-specific information disparity.
While hypernetworks offer an environment to generate model weights for portions of the network that lack client-sourced parameters, their naive use results in significant computational and memory overhead. We introduce an LRF-based compression scheme to optimize hypernetwork operations, significantly reducing resource demands and enabling practical deployment of hypernetworks in federated learning.
@@ 2. Related work and Background
We first positioning our work within current literature and discuss foundational concepts and methodologies relevant to our study.
Device Heterogeneity in Federated Learning. While federated learning research has mostly focused on homogeneous models with little consideration of client-side computing capabilities (Diao et al., 2020), supporting federated learning for heterogeneous clients is crucial. It allows for tailored models that adapt to diverse computational capabilities, data distributions, and privacy requirements of various devices (Zhang et al., 2022; Bonawitz et al., 2019; Li et al., 2020a). Heterogeneous federated learning leverages each client’s unique capabilities, ensuring that even resource-constrained devices contribute effectively to enhance model robustness, maximize the potential of collective data, and lead to better generalization across tasks and environments (Lin et al., 2020).
One early approach to achieve device heterogeneity in federated learning is federated distillation (FD) (Li and Wang, 2019; Seo et al., 2022), which leverages knowledge distillation (Hinton et al., 2015). Instead of sharing model weights, FD aggregates class scores from a large public dataset computed locally at each client to accommodate diverse client models. However, FD requires the server to maintain access to a well-balanced public shared dataset, which can be impractical in privacy-sensitive settings. Furthermore, the availability of such datasets may also introduce biases, limiting model the generalizability across diverse client data distributions.
An alternative approach involves pruning local models based on the Lottery Ticket hypothesis (Frankle and Carbin, 2018). In LotteryFL (Li et al., 2020b) and Hermes (Li et al., 2021a), each client learns a subnetwork of the global model (i.e., lottery ticket network) through iterative pruning, rewinding, and training, exchanging only sparse pruned parameters. However, this iterative process is practically challenging for resource-limited clients, as they must initially train the heavy full model to determine which weights to prune. This requires a large number of rounds for model convergence and can cause training delays. Furthermore, aggressively pruned models may lose significant amounts of information, affecting both model convergence and final accuracy.
@@
Multi-Exit Network Architecture. As the name implies, multi-exit architectures, inject multiple (early) exit points within a neural network to enhance inference latency by enabling adaptive inference and eliminating unnecessary operations (Laskaridis et al., 2021; Lee et al., 2024a). Teerapittayanon et al. introduced BranchNet, a pioneering multi-exit deep neural network featuring multiple classifiers integrated within the baseline model to improve computational latency (Teerapittayanon et al., 2016). Given that mobile and embedded devices often experience lagging inference due to resource limitations, multi-exit architectures have been often applied in various real-world sensing applications within mobile and embedded computing environments (Kouris et al., 2022; Leontiadis et al., 2021; Kim et al., 2020). However, while these existing methods can improve performance during inference, they hold challenges in local training, as all exits must be trained. As a result, heterogeneous clients cannot benefit from this approach for the training phase.
Hypernetworks. Hypernetworks are neural networks designed to generate weights for another neural network (Ha et al., 2016), offering flexibility in applications like computer vision (Jia et al., 2016; Klocek et al., 2019) and 3D scene representation (Littwin and Wolf, 2019; Sitzmann et al., 2020). In federated learning, hypernetworks are used to enhance model initialization and personalization (Shamsian et al., 2021; Litany et al., 2022) and to determine aggregation ratios for model personalization (Ma et al., 2022). However, in existing federated learning using hypernetworks, the training of the hypernetwork takes place on each local client, imposing an additional computational burden on the clients.
@@ 3.2. Preliminary Study
In Figure 3 (a) we also plot four additional cases where our proposed solutions in Section 3.1 were applied in different combinations. As we allocate clients with their largest-adoptable model, we text cases where neither the multi-exit approach nor the hypernetwork is used (“ME X/HN X”), cases where one of the two is used (“ME O/HN X” and “ME X/HN O”), and a final case there both approaches are integrated into the federated learning process (“ME O/HN O”). We use these plots to examine the feasibility of applying our solutions for federated learning scenarios. The results in Figure 3 (a) suggest that in the absence of the multi-exit architecture, convergence issues arise due to misaligned feature spaces, while the system faces issues due to per-layer information disparity when without the hypernetwork-based weight generations. In particular, without a multi-exit architecture, clients with limited resources would provide updates that do not align with those from clients running deeper models; thus, hindering global convergence. Similarly, without hypernetworks, the server cannot effectively gather (enough) weights for deeper layers, leaving gaps in layer-specific knowledge aggregation. As a result, the case with both multi-exit and hypernetwork (“ME O/HN O”) shows the highest performance, close to that of an ideal case where all clients exploit the full model (“Large (All)”). These results emphasize the need for integrating both the multi-exit architecture and hypernetwork approaches to ensure robust federated learning across heterogeneous clients.
@@
Without multi-exit layers, the one-layer (S) and two-layer (M) models differ significantly from the full model (L), with CKA values of 0.6761 and 0.6764, respectively. However, incorporating multi-exit architecture improves feature alignment noticeably, yielding CKA values of 0.7812 and 0.9407 for these configurations. This improved alignment underlines how multi-exit architectures help to preserve feature consistency across clients with varying model depths. Comparisons like “M-L (X-O)” and “M-L (O-X)” in Figure 3 (b) show considerable differences in feature alignment depending on the presence of multi-exit layers, with CKA values of 0.7165 and 0.8343, respectively. The improvement in CKA values implies the critical role of multi-exit networks in not only balancing the computational load, but also in maintaining coherent feature spaces. These findings highlight the effectiveness of multi-exit architectures in depth-wise split networks for federated learning, enhancing feature space alignment and potentially boosting overall model performance.
@@
Overall, our preliminary results suggest the potential effectiveness of leveraging the multi-exit architecture and hypernetworks for heterogeneous federated learning. The combination of multi-exit architecture for feature alignment and hypernetworks for filling in missing layers provides a synergistic solution to the challenges posed by client heterogeneity. Using these results as empirical evidence, we discuss the details of our proposed HypeMeFed design in the following section.
@@ 4.1. Overview of HypeMeFed
The server distributes global model parameters to clients based on their computing resources. Here, we note that the server can easily acquire information about the clients’ computing resources via a simple request. The full global model is configured with multiple exit layers organized in a depth-wise manner ( ). Clients with higher computational resources receive the full model, while the ones that are limited receive a smaller subset of the model split at earlier exit layers ( ). (c.f., Sec. 4.2).
@@ 4.2. Multi-Exit Neural Network Architecture
The placement and number of exit layers in HypeMeFed are key design choices. In this work, we use three exit layers positioned at one-third, two-thirds, and the end of the federated learning model to exploit a range of features that the model embeds (Park et al., 2023). This setup allows the categorization of clients into three groups based on their computational capabilities: the first group, with limited resources, uses a subnetwork up to the first exit layer; the second group, with moderate resources, uses up to the second exit; and the third group, with ample resources, uses the full network with all three exits. During local training, clients compute outputs for all exit layers included in their allocated model. For instance, a client with sufficient resources generates predictions from all three exit layers and a resource limited client will compute only up to the first. Each layer’s output contributes to the overall loss, summing the losses from all exits—a standard practice in multi-exit neural networks (Laskaridis et al., 2021; Teerapittayanon et al., 2016; Lee et al., 2024a). In this work, we adopt a simple categorization of clients’ resource heterogeneity in three levels and accordingly use three exits. However, depending on design choices, the granularity of client resource heterogeneity can adjusted. As we explain later, HypeMeFed is a general framework that does not depend on the number of exits; thus, the number of exits is not a constraint. Additionally, the optimal placement of exit layers can vary based on the model architecture and the characteristics of the data, and can be determined experimentally by the system designers. In this paper, rather than searching for the optimal exit positions for each dataset, we have assigned the exit positions to uniformly split the model, dividing the model into thirds, to validate a general configuration.
@@
HypeMeFed facilitates seamless aggregation of trained model parameters despite heterogeneous network architectures by training and sharing all exit layer parameters. Specifically, HypeMeFed can easily be integrated with various federated learning schemes like FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020a), and SCAFFOLD (Karimireddy et al., 2020) without modifying the local training and aggregation process. We emphasize that this flexibility further enables HypeMeFed to effectively address the prevalent statistical data heterogeneity observed in real-world usage scenarios. In this work, HypeMeFed exploits the FedAvg approach, a straightforward method for averaging model parameters during server-side model aggregation.
@@ 4.3. Efficient Hypernetwork Design for Multi-exit Architectures
To integrate hypernetworks in federated learning, a straightforward approach involves vectorizing parameters from preceding layers and feeding them into a multi-layer perceptron (MLP) model. The central server, which aggregates parameters from clients with different computational capabilities, possesses information on both preceding and subsequent layer parameters. This allows the server to utilize parameter information from clients computing up to the second and third exits to train the hypernetwork and understand relationships between adjacent exit parameters. However, neural networks typically entail numerous parameters, demanding substantial memory and time for hypernetwork training. Despite HypeMeFed’s hypernetworks operating on a resource-rich server (a practical configuration for real-world federated learning systems), a naïve hypernetwork design can be computationally intensive and memory-consuming, posing practical challenges in real-world deployments, which often makes them impractical with respect to the training time and memory usage (c.f., Section 5.4). To tackle this, we propose a more efficient hypernetwork architecture that balances effective parameter generation with the computational constraints of federated learning environments.
HypeMeFed introduces an efficient hypernetwork architecture using a low-rank factorization (LRF) strategy. Instead of vectorizing and reconstructing all model parameters, which is a computationally intensive approach, LRF compresses essential neural network parameters to reduce the computational load and memory requirements. By leveraging the inherent sparsity in neural networks, similar to what is leveraged in pruning techniques, this approach is both feasible and efficient (He et al., 2017; Frankle and Carbin, 2018; Evci et al., 2022).
@@ 4.4. Hypernetwork Implementations
Given the early-exit architecture of HypeMeFed, which in this paper splits the model into three subnetworks, we employ two hypernetworks: one to predict the second exit parameters from the first exit parameters () and another to predict the third exit parameters from the second (). These hypernetworks are trained at the server using parameter samples collected from clients each round. Clients executing the model up to the second exit provide samples for , while those executing the full model provide samples for . A briefly mentioned earlier, it is important to note that HypeMeFed is not limited to a specific number of exits. Our scheme can be generalized to any number of exits, with hypernetworks employed in an auto-regressive manner, where the hypernetwork for each exit () predicts the parameters of the next exit based on the current one. This scalable design allows HypeMeFed to accommodate various levels of model depth and client capabilities, making it adaptable to different deployment scenarios as discussed in Section 5.4.
@@
HypeMeFed selectively generates only the core parameters of the feature extractor, specifically the convolutional layers, while excluding the batch normalization (BN) and classification layers. This approach is based on two key empirical observations: (i) changes to the BN and classification layers do not significantly enhance global model performance, and (ii) since these layers are closely related to local data, globally generating them can dilute the unique knowledge generated at each client, compromising model personalization. We demonstrate this in Section 5.3, where we evaluate personalization accuracy after a few epochs of local fine-tuning of the global model with client data.
Once client model parameter data is collected at each federated learning round and the hypernetworks are trained using the operations described above, HypeMeFed uses the fine-tuned hypernetworks to predict subsequent layer parameters based on the compressed representations. This approach ensures that clients with varying computational capacities can fairly participate in the federated learning process by generating necessary weight information for effective model convergence and improved performance.